# Enabling Automatic Differentiation with Mollified Graph Neural Operators

### v1

@@ Abstract
Physics-informed neural operators offer a powerful framework for learning solution operators of partial differential equations (PDEs) by combining data and physics losses. However, these physics losses rely on derivatives. Computing these derivatives remains challenging, with spectral and finite difference methods introducing approximation errors due to finite resolution. Here, we propose the mollified graph neural operator ($m$GNO), the first method to leverage automatic differentiation and compute exact gradients on arbitrary geometries. This enhancement enables efficient training on irregular grids and varying geometries while allowing seamless evaluation of physics losses at randomly sampled points for improved generalization. For a PDE example on regular grids, $m$GNO paired with autograd reduced the L2 relative data error by 20× compared to finite differences, although training was slower. It can also solve PDEs on unstructured point clouds seamlessly, using physics losses only, at resolutions vastly lower than those needed for finite differences to be accurate enough. On these unstructured point clouds, $m$GNO leads to errors that are consistently 2 orders of magnitude lower than machine learning baselines (Meta-PDE) for comparable runtimes, and also delivers speedups from 1 to 3 orders of magnitude compared to the numerical solver for similar accuracy. $m$GNOs can also be used to solve inverse design and shape optimization problems on complex geometries.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Mollified GNO kernel’s neighborhood and weights versus those of the vanilla GNO kernel with differing point densities.
@@
PDEs are critical for modeling physical phenomena relevant for scientific applications. Unfortunately, numerical solvers become very expensive computationally when used to simulate large-scale systems. To avoid these limitations, neural operators, a machine learning paradigm, have been proposed to learn solution operators of PDEs (Azizzadenesheli et al., 2024). Neural operators learn mappings between function spaces rather than finite-dimensional vector spaces, and approximate solution operators of PDE families (Li et al., 2020b; Kovachki et al., 2023). One example is the Fourier Neural Operator (FNO) (Li et al., 2020a), which relies on Fourier integral transforms with kernels parameterized by neural networks. Another example, the Graph Neural Operator (GNO) (Li et al., 2020b), implements kernel integration with graph structures and is applicable to complex geometries and irregular grids. The GNO has been combined with FNOs in the Geometry-Informed Neural Operator (GINO) (Li et al., 2023) to handle arbitrary geometries when solving PDEs. Neural operators have been successfully used to solve PDE problems with significant speedups (Li et al., 2021a; Kurth et al., 2022). They have shown great promise, primarily due to their ability to receive input functions at arbitrary discretizations and query output functions at arbitrary points, and also due to their universal operator approximation property (Kovachki et al., 2021).
Figure 2: Qualitative assessment of the smoothness of derivatives. We fit GINO with a PINO loss to the differentiable function $u(x,y)=\sin(4\pi xy)$ on $[0,1]^{2}$, and compare the analytic ground truth derivative $\partial_{xx}u(x,y)=-16\pi^{2}y^{2}\sin(4\pi xy)$ (left) with the numerical derivatives obtained using automatic differentiation (middle) and finite differences (right).
@@
However, purely data-driven approaches may underperform in situations with limited or low resolution data (Li et al., 2021b), and may be supplemented using knowledge of physics laws (Karniadakis et al., 2021) as additional loss terms, as done in Physics-Informed Neural Networks (PINNs) (Raissi et al., 2017a, b, 2019). In the context of neural operators, the Physics-Informed Neural Operator (PINO) (Li et al., 2021b) combines training data (when available) with a PDE loss at a higher resolution, and can be finetuned on a given PDE instance using only the equation loss to provide a nearly-zero PDE error at all resolutions.
A major challenge when using physics losses is to efficiently compute derivatives without sacrificing accuracy, since numerical errors on the derivatives are compounded in the physics losses and output solutions. Approximate derivatives can be computed using finite differences (FDs), but can require a very high resolution grid to be sufficiently accurate, thus becoming intractable for fast-varying dynamics. Numerical derivatives can also be computed using Fourier differentiation, but this requires smoothness, uniform grids, and performs best when applied to periodic problems. In contrast, automatic differentiation computes exact derivatives using repeated applications of the chain rule and scales better to large-scale problems and fast-varying dynamics. Unlike numerical differentiation methods which introduce errors, automatic differentiation gives exact gradients, ensuring the accuracy required for physics losses, making it the preferred approach for physics-informed machine learning.
Approach.
 We propose a fully differentiable modification to GNOs to allow for the use automatic differentiation when computing derivatives and physics losses, which was previously prohibited by the GNO’s non-differentiability. More precisely, we replace the non-differentiable indicator function in the GNO kernel integration by a differentiable weighted function. This is inspired by mollifiers in functional analysis (Evans, 2010), which are used to approximate, regularize, or smooth functions. As a result, our differentiable mollified GNO ($m$GNO) is the first method capable of computing exact derivatives at arbitrary query points. The resulting $m$GNO can, in particular, be used within GINO to learn efficiently and accurately the solution operator of families of large-scale PDEs with varying geometries without data using physics losses.
We test the proposed approach on Burgers’ equation with regular grids, and nonlinear Poisson and hyperelasticity equations with varying domain geometries. Figure 3 displays examples of solutions, highlighting the complexity of the geometries considered). We show that physics losses are often critical, highlighting the need for efficient and accurate methods to compute derivatives. Using autograd instead of FDs leads to 20× reductions of the L2 relative data loss for Burgers’ equation on regular grids, suggesting that the Autograd physics loss better captures the physics underlying the data. Autograd $m$GINO performs seamlessly for the PDEs on unstructured point clouds, while FDs are not sufficiently accurate at the same training resolution and would need at least 9× more points to compute reasonable derivatives. Autograd $m$GINO achieves a relative error 2-3 orders of magnitude lower than the machine learning baselines (Meta-PDE (Qin et al., 2022) with LEAP and MAML) considered for a comparable running time, and enjoys speedups of 20-25× and 3000-4000× compared to the numerical solver for similar accuracy on the Poisson and hyperelasticity equations. Furthermore, as a result of its differentiability, $m$GINO can be used seamlessly for solving inverse design and shape optimization problems on complex geometries, as demonstrated with an airfoil design problem.
@@ 2.1 Neural Operators
Neural operators compose linear integral operators $\mathcal{K}$ with pointwise non-linear activation functions $\sigma$ to approximate non-linear operators. A neural operator is defined as
@@
where $\mathcal{Q}$ and $\mathcal{P}$ are the pointwise neural networks that encode (lift) the lower dimension function onto a higher dimensional space and project it back to the original space, respectively. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters of a FNO model consist of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. (2024) maintain a comprehensive open-source library for learning neural operators in PyTorch, which serves as the foundation for our implementation.
@@
where $R_{\phi}$ is the Fourier transform of a periodic function $\kappa$ parameterized by $\phi$. On a uniform mesh, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform. The FNO architecture is displayed in Figure 8.
@@
 GNO (Li et al., 2020b) implements kernel integration with graph structures and is applicable to complex geometries and irregular grids. The GNO kernel integration shares similarities with the message-passing implementation of graph neural networks (GNN) (Battaglia et al., 2016). However, GNO defines the graph connection in a ball in physical space, while GNN assumes a fixed set of neighbors. The GNN nearest-neighbor connectivity violates discretization convergence and degenerates into a pointwise operator at high resolutions, leading to a poor approximation of the operator. In contrast, GNO adapts the graph based on points within a physical space, allowing for universal approximation of operators.
@@
 GINO (Li et al., 2023) proposes to combine a FNO with GNOs to handle arbitrary geometries. More precisely, the input is passed through three main neural operators,
@@
First, a GNO encodes the input given on an arbitrary geometry into a latent space with a regular geometry. The encoded input can be concatenated with a signed distance function evaluated on the same grid if available. Then, a FNO is used as a mapping on that latent space for global integration. Finally, a GNO decodes the output of the FNO by projecting that latent space representation to the output geometry. The GINO architecture is displayed in Figure 9.
@@ 2.2 Physics-Informed Machine Learning
 In the context of solving PDEs, a PINN (Raissi et al., 2017a, b, 2019) is a neural network representation of the solution of a PDE, whose parameters are learned by minimizing the distance to the reference PDE solution and deviations from known physics laws such as conservation laws, symmetries and structural properties, and from the governing differential equations. PINNs minimize a composite loss
@@
where $\mathcal{L}_{\text{data}}$ measures the error between data and model predictions wile $\mathcal{L}_{\text{physics}}$ penalizes deviations away from physics laws. PINN overcomes the need to choose a discretization grid of most numerical solvers, but only learns the solution for a single PDE instance and do not generalize to other instances without further optimization. Many modified versions of PINNs have also been proposed and used to solve PDEs in numerous contexts (Jagtap et al., 2020; Cai et al., 2022; Yu et al., 2022). When data is not available, we can try to learn the solution by minimizing $\mathcal{L}_{\text{physics}}$ only.
While physics losses can prove very useful, the resulting optimization task can be challenging and prone to numerical issues. The training loss typically has worse conditioning as it involves differential operators that can be ill-conditioned. In particular, Krishnapriyan et al. (2021) showed that the loss landscape becomes increasingly complex and harder to optimize as the physics loss coefficient $\lambda$ increases. The model could also converge to a trivial or non-desired solution that satisfies the physics laws on the set of points where the physics loss is computed (Leiteritz & Pflüger, 2021). There could also be conflicts between the multiple loss terms. Even without such conflicts, the losses can vary significantly in magnitude, leading to unbalanced back-propagated gradients during training (Wang et al., 2021). Manually tuning the loss coefficients can be computationally expensive, especially as the number of loss terms increases. Note that strategies have been developed to adaptively update the loss coefficients and mitigate this issue (Chen et al., 2018; Heydari et al., 2019; Bischof & Kraus, 2021).
@@
 In PINO (Li et al., 2021b), a FNO is trained with training data (when available) and physics losses at a higher resolution, allowing for near-perfect approximations of PDE solution operators. To further improve accuracy at test time, the trained model can be finetuned on the given PDE instance using only the equation loss and provides a nearly-zero error for that instance at all resolutions. PINO has been successfully applied to many PDEs (Song et al., 2022; Meng et al., 2023; Rosofsky et al., 2023). In practice, the PDE loss vastly improves generalization, physical validity, and data efficiency in operator learning compared to purely data-driven methods.
@@
 Various approaches enforce physics laws in surrogate models without using physics losses. This can be achieved, for instance, by using projection layers (Jiang et al., 2020; Duruisseaux et al., 2024; Harder et al., 2024), by finding optimal linear combinations of learned basis functions that solve a PDE-constrained optimization problem (Negiar et al., 2022; Chalapathi et al., 2024), by leveraging known characterizations and properties of the solution operator as for divergence-free flows (Richter-Powell et al., 2022; Mohan et al., 2023; Xing et al., 2024), or Hamiltonian systems with their symplectic structure (Burby et al., 2020; Jin et al., 2020; Chen & Tao, 2021; Duruisseaux et al., 2023).

### v2

@@ Abstract
Physics-informed neural operators offer a powerful framework for learning solution operators of partial differential equations (PDEs) by combining data and physics losses. However, these physics losses require the efficient and accurate computation of derivatives. Computing these derivatives remains challenging, with spectral and finite difference methods introducing approximation errors due to finite resolution. Here, we propose the mollified graph neural operator (mmGNO), the first method to leverage automatic differentiation and compute exact gradients on arbitrary geometries. This enhancement enables efficient training on arbitrary point clouds and irregular grids with varying geometries while allowing the seamless evaluation of physics losses at randomly sampled points for improved generalization. For a PDE example on regular grids, mmGNO paired with Autograd reduced the L2 relative data error by 20× compared to finite differences, suggesting it better captures the physics underlying the data. It can also solve PDEs on unstructured point clouds seamlessly, using physics losses only, at resolutions vastly lower than those needed for finite differences to be accurate enough. On these unstructured point clouds, mmGNO leads to errors that are consistently 2 orders of magnitude lower than machine learning baselines (Meta-PDE, which accelerates PINNs) for comparable runtimes, and also delivers speedups from 1 to 3 orders of magnitude compared to the numerical solver for similar accuracy. mmGNOs can also be used to solve inverse design and shape optimization problems on complex geometries.
@@ 1 Introduction
Figure 1: (left) Mollified GNO kernel’s neighborhood and weights versus those of the vanilla GNO kernel with differing point densities. (right) Examples of weight functions (equation 9 - equation 12) for mmGNO.
@@
PDEs are critical for modeling physical phenomena relevant to scientific applications. Unfortunately, numerical solvers become very expensive computationally when used to simulate large-scale systems. To avoid these limitations, neural operators, a machine learning paradigm, have been proposed to learn solution operators of PDEs (Azizzadenesheli et al., 2024; Berner et al., 2025). Neural operators learn mappings between function spaces rather than finite-dimensional vector spaces, and, in particular can be used to approximate solution operators of PDE families (Li et al., 2020b; Kovachki et al., 2023). One example is the Fourier Neural Operator (FNO) (Li et al., 2020a), which relies on Fourier integral transforms with kernels parameterized by neural networks. Another example, the Graph Neural Operator (GNO) (Li et al., 2020b), implements kernel integration with graph structures and is applicable to complex geometries and irregular grids. The GNO has been combined with FNOs in the Geometry-Informed Neural Operator (GINO) (Li et al., 2023) to handle arbitrary geometries when solving PDEs. Neural operators have been successfully used to solve PDE problems with significant speedups (Kurth et al., 2022; Gopakumar et al., 2023; Wen et al., 2023; Ghafourpour et al., 2025). They have shown great promise, primarily due to their ability to receive input functions at arbitrary discretizations and query output functions at arbitrary points, and also due to their universal operator approximation property (Kovachki et al., 2021).
However, purely data-driven approaches may underperform in situations with limited or low-resolution data (Li et al., 2021b), and may be supplemented using knowledge of physics laws (Karniadakis et al., 2021) as additional loss terms, as done in Physics-Informed Neural Networks (PINNs) (Raissi et al., 2017a; b; 2019). In the context of neural operators, the Physics-Informed Neural Operator (PINO) (Li et al., 2021b) combines training data (when available) with a PDE loss at a higher resolution, and can be fine-tuned on a given PDE instance using only the equation loss to provide a nearly zero PDE error at all resolutions.
A major challenge when using physics losses is to efficiently compute derivatives without sacrificing accuracy, since numerical errors on the derivatives are compounded in the physics losses. Approximate derivatives can be computed using finite differences, but can require a very high resolution grid to be sufficiently accurate, thus becoming intractable for fast-varying dynamics. Numerical derivatives can also be computed using Fourier differentiation, but this requires smoothness, uniform grids, and performs best when applied to periodic problems. In contrast, automatic differentiation computes exact derivatives using repeated applications of the chain rule and scales better to large-scale problems and fast-varying dynamics. Unlike numerical differentiation methods, which introduce errors, automatic differentiation gives exact gradients, ensuring the accuracy required for physics losses, making it the preferred approach for physics-informed machine learning.
Approach.
We propose a fully differentiable modification to GNOs to allow for the use of automatic differentiation when computing derivatives and physics losses, which was previously prohibited by the GNO’s non-differentiability. More precisely, we replace the non-differentiable indicator function in the GNO kernel integration by a differentiable weighting function. This is inspired by mollifiers in functional analysis (Evans, 2010), which are used to approximate, regularize, or smooth functions.
As a result, our differentiable mollified GNO (mmGNO) is the first method capable of computing exact derivatives at arbitrary query points. The resulting mmGNO can, in particular, be used within GINO to learn efficiently and accurately the solution operator of families of large-scale PDEs with varying geometries without data using physics losses. A data loss can help further improve the results when reference simulations are available, even if these are of very low resolution, when paired with a high-resolution physics loss.
@@
Summary of Results.
We test our approach on Burgers’ and Navier–-Stokes equations with regular grids, and on nonlinear Poisson and hyperelasticity equations with varying domain geometries. Figure˜2 shows example solutions, highlighting the complexity of the geometries considered. Physics losses prove critical, emphasizing the need for accurate and efficient derivative computation.
Using Autograd instead of finite differences reduces the relative L2 data loss by 20× for Burgers’ equation on regular grids, showing that the Autograd physics loss captures the underlying physics more accurately. In hybrid training with noisy reference data, the Autograd-based $m\mathrm{GNO}\circ\mathrm{FNO}$ remains robust and accurate, maintaining low PDE residuals and data loss.
It also excels at learning the lid cavity flow example governed by the Navier–Stokes equations. For this problem, access to data is essential, but a small number of randomly sampled points (with Autograd) at each iteration achieves excellent results when trained with a hybrid loss. Including a physics loss is also critical, as it ensures PDE fidelity and guides the network toward physically consistent solutions even at low resolution.
Autograd mmGINO performs seamlessly on unstructured point clouds, while finite difference derivatives are not sufficiently accurate at the same training resolution and would need at least 9× more points. Autograd mmGINO achieves a relative error 2-3 orders of magnitude lower than the machine learning baselines (Meta-PDE (Qin et al., 2022) using LEAP and MAML) considered for a comparable running time, and enjoys speedups of 20-25× and 3000-4000× compared to the numerical solver for similar accuracy on the Poisson and hyperelasticity equations. Furthermore, as a result of its differentiability, mmGINO can be used seamlessly for solving inverse design and shape optimization problems on complex geometries, as demonstrated with an airfoil design problem.
@@ 2.1 Neural Operators
Neural operators compose linear integral operators $\mathcal{K}$ with pointwise nonlinear activation functions $\sigma$ to approximate nonlinear operators. A neural operator is defined as
@@
where $\mathcal{Q}$ and $\mathcal{P}$ are the pointwise neural networks that encode (lift) the lower dimension function onto a higher-dimensional space and project it back to the original space, respectively. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions.
The parameters of a neural operator consist of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. (2025) maintains a comprehensive open-source library for learning neural operators in PyTorch, which serves as the foundation for our implementation.
@@
where $R_{\phi}$ is the Fourier transform of a periodic function $\kappa$ parameterized by $\phi$. When the discretizations of both the input and output functions are given on a uniform mesh, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform. We note that the FNO architecture is fully differentiable. A depiction of the FNO architecture is provided in Figure˜14.
@@
A GNO (Li et al., 2020b) implements kernel integration with graph structures and is applicable to complex geometries and irregular grids. The GNO kernel integration shares similarities with the message-passing implementation of graph neural networks (GNNs) (Battaglia et al., 2016). However, GNO defines the graph connection in a ball in continuous physical space, while GNN assumes a fixed set of neighbors in a discrete graph. The GNN nearest-neighbor connectivity violates discretization convergence and degenerates into a pointwise operator at high resolutions, leading to a poor approximation of the operator. In contrast, GNO adapts the graph based on points within a physical space, allowing for universal approximation of operators.
@@
GINO (Li et al., 2023) proposes to combine an FNO with GNOs to handle arbitrary geometries. More precisely, the input is passed through three main neural operators,
@@
First, a GNO encodes the input given on an arbitrary geometry into a latent space with a regular geometry. The encoded input can be concatenated with a signed distance function evaluated on the same grid if available. Then, an FNO is used as a mapping on that latent space for efficient global integration. Finally, a GNO decodes the output of the FNO by projecting that latent space representation to the output geometry. As a result, GINO can represent mapping from one complex geometry to another. A depiction of the GINO architecture is displayed in Figure˜15.
@@ 2.2 Physics-Informed Machine Learning
In the context of solving PDEs, a PINN (Raissi et al., 2017a; b; 2019) is a neural network representation of the solution of a PDE, whose parameters are learned by minimizing the distance to the reference PDE solution and deviations from known physics laws such as conservation laws, symmetries and structural properties, and from the governing differential equations. PINNs minimize a composite loss
@@
where $\mathcal{L}_{\text{data}}$ measures the error between data and model predictions while $\mathcal{L}_{\text{physics}}$ penalizes deviations away from physics laws. PINN overcomes the need to choose a discretization grid of most numerical solvers, but only learns the solution for a single PDE instance and does not generalize to other instances without further optimization. Many modified versions of PINNs have also been proposed and used to solve PDEs in numerous contexts (Jagtap et al., 2020; Cai et al., 2022; Yu et al., 2022). When data is not available, we can try to learn the solution by minimizing $\mathcal{L}_{\text{physics}}$ only.
While physics losses can prove very useful, the resulting optimization task can be challenging and prone to numerical issues. The training loss typically has poor conditioning as it involves differential operators that can be ill-conditioned. In particular, Krishnapriyan et al. (2021) showed that the loss landscape becomes increasingly complex and harder to optimize as the physics loss coefficient $\lambda$ increases. The model could also converge to a trivial or non-desired solution that satisfies the physics laws on the set of points where the physics loss is computed (Leiteritz & Pflüger, 2021). There could also be conflicts between the multiple loss terms when their gradients point in opposite directions. Even without such conflicts, the losses can vary significantly in magnitude, leading to unbalanced backpropagated gradients during training (Wang et al., 2021), and thus to the contribution and reduction of certain losses being relatively negligible during training. Manually tuning the loss coefficients can be computationally expensive, especially as the number of loss terms increases. Note that strategies have been developed to adaptively update the loss coefficients and mitigate this issue (Chen et al., 2018; Heydari et al., 2019; Bischof & Kraus, 2021).
@@
In the physics-informed neural operator approach (Li et al., 2021b), an FNO is trained with low-resolution simulations (or even without data) and high-resolution physics losses, allowing for near-perfect approximations of PDE solution operators. To further improve accuracy at test time, trained models can be fine-tuned on a given PDE instance using only the equation loss and provide a near-zero error at all resolutions. PINO has been successfully applied to many PDEs (Song et al., 2022; Meng et al., 2023; Rosofsky et al., 2023). In practice, the PDE loss vastly improves generalization, physical validity, and data efficiency in operator learning compared to purely data-driven methods.
@@
Various approaches enforce physics laws in surrogate models without using physics losses. This can be achieved, for instance, by using projection layers (Jiang et al., 2020; Duruisseaux et al., 2024; Harder et al., 2024), by finding optimal linear combinations of learned basis functions that solve a PDE-constrained optimization problem (Negiar et al., 2022; Chalapathi et al., 2024), by leveraging known characterizations and properties of the solution operator as for divergence-free flows (Richter-Powell et al., 2022; Mohan et al., 2023; Xing et al., 2024), or Hamiltonian systems with their symplectic structure (Burby et al., 2020; Jin et al., 2020; Chen & Tao, 2021; Duruisseaux et al., 2023). These approaches are particularly advantageous when the system is well-understood and physical constraints can be explicitly encoded, ensuring strict compliance with the underlying physics. They also tend to require less computational overhead than methods that rely on minimizing physics losses. However, such approaches are often tailored to specific PDE structures and are limited to dynamical systems with well-characterized solutions. In contrast, physics-loss-based methods offer greater flexibility and broader applicability across a wide range of PDEs, including systems whose solution properties are not fully understood. By incorporating physics-based loss terms during training, as in PINNs and PINOs, the governing equations serve as regularizers, offering greater flexibility, broad applicability across diverse PDEs, and the advantage of requiring only knowledge of the underlying equations. Hybrid strategies are also possible, in which certain constraints are enforced directly in the architecture while others are imposed through physics losses during training. For instance, in incompressible Navier–Stokes problems, a physics loss can penalize deviations from the momentum equation, while the divergence-free condition can be enforced explicitly in the model (e.g. using projections as done in Jiang et al. (2020) and Duruisseaux et al. (2024)).
Physics-Infused Transformer Architectures.
Recent parallel research has explored embedding certain structures of PDEs into Transformer architectures. Transolver (Wu et al., 2024; Luo et al., 2025) introduces a physics-attention mechanism that adaptively groups mesh points with similar physical states, enabling the model to capture complex correlations across arbitrary geometries efficiently. Unisolver (Zhou et al., 2025) conditions a Transformer on complete physics information, incorporating both domain-wise components, such as equation symbols and coefficients, and point-wise components such as boundary conditions, achieving broad applicability across diverse PDEs. PDEFormer (Ye et al., 2025) represents PDEs as computational graphs and combines symbolic and numerical information through a graph Transformer and implicit neural representation, allowing mesh-free solutions and efficient handling of varied PDE types. Collectively, these architectures illustrate an alternative approach to integrating physical knowledge into Transformer structures, offering complementary strategies for improving the accuracy and flexibility of machine learning surrogates for PDE solvers.